OpenStudy (anonymous):

Solve –2(x – 1) = 12

OpenStudy (anonymous):

sorry.. -5

OpenStudy (anonymous):

-2x+2=12 -2x=10 x=-5

OpenStudy (anonymous):

@toadytica305 :Plz dont give the final answer ,that against the C.O.C ...Just explain from where u got the answer and let him solve it ....

OpenStudy (anonymous):

@Eyad You just gave him/her the answer as well..

OpenStudy (anonymous):

@toadytica305 :I gave him the answer by explaining from where i got it ,not the final answer as u did !